# Radio Map-Based Spectrum Sharing for Joint Communication and Sensing

### v1

@@ II-B CHANNEL MODEL
In this subsection, we introduce the channel model. We assume that BSs and radars are equipped with a single antenna. This is because, in MIMO systems, the precoding scheme that exploits the multiplexing gain requires perfect and instantaneous CSI [34]. With only large-scale CSI available, multiple antennas are equivalent to a combined single antenna for both BSs and radars. The interfered user is assumed to have $N_{c}$ antennas, which is used to receive different signals from distributed BSs. On this basis, we denote the BS-user link and interference links, i.e., BS-radar link and radar-user link as $\mathbf{H}_{c}\in\mathbb{C}^{M_{c}\times N_{c}}$, $\mathbf{h}_{c\rightarrow r_{i}}\in\mathbb{C}^{M_{c}\times 1}$, and $\mathbf{h}_{r_{i}\rightarrow c}\in\mathbb{C}^{1\times M_{r}}$, respectively, where $i$ denotes the index number of radars. They are modeled by the composite channel model:
@@
where $\mathbf{L}_{c}\in\mathbb{C}^{M_{c}\times M_{c}}$, $\mathbf{L}_{c\rightarrow r_{i}}\in\mathbb{C}^{M_{c}\times M_{c}}$, and $l_{r_{i}\rightarrow c}$ denote the large-scale fading. $\mathbf{L}_{c}$ and $\mathbf{L}_{c\rightarrow r_{i}}$ are diagonal matrices whose $j$th diagonal elements are denoted as $l_{c_{j}}$ and $l_{c_{j}\rightarrow r_{i}}$. $\mathbf{S}_{c}\in\mathbb{C}^{M_{c}\times N_{c}}$, $\mathbf{s}_{c\rightarrow r_{i}}\in\mathbb{C}^{M_{c}\times 1}$, and $\mathbf{s}_{r_{i}\rightarrow c}\in\mathbb{C}^{1\times N_{c}}$ denote the small-scale fading, whose elements are uniformly and independently conformed to the Gaussian distribution, i.e., $\mathcal{CN}(0,1)$. For the two-way links between the radars and target, i.e., $\mathbf{H}_{r}=[\mathbf{h}_{r_{1}},....,\mathbf{h}_{r_{M_{r}}}]\in\mathbb{C}^{M_{r}\times M_{r}}$, they are modeled by the determined channel model. This is because, in the case of surveillance radars, the detected target is more likely to be in the sky. The corresponding air-to-ground channel is dominated by the line-of-sight path with little random fading.
@@ II-C SIGNAL MODEL
As for the sensing system, radars use orthogonally-coded signals and independently detect a mutual target. Assuming that the delay and Doppler shift are perfectly compensated, the received echo of radar $i$ is given by
$y_{r_{i}}(t)=[\mathbf{h}_{r_{i}}(t)]^{H}\mathbf{x}_{r}(t)+[\mathbf{h}_{c\rightarrow r_{i}}(t)]^{H}\mathbf{x}_{c}(t)+w_{r_{i}}(t),$ (2)
where $\mathbf{x}_{r}(t)=[x_{r_{1}}(t),...,x_{r_{M_{r}}}(t)]^{T}\in\mathbb{C}^{M_{r}\times 1}$ and $\mathbf{x}_{c}(t)=[x_{c_{1}}(t),...,x_{c_{M_{c}}}(t)]^{T}\in\mathbb{C}^{M_{c}\times 1}$ denote the transmit signals of radars and BSs, and $w_{r_{i}}(t)$ is the additive white Gaussian noise. To distinguish the echo of radar $i$ from that of others, the radar receiver is equipped with a matched filter (MF), i.e., $\frac{x_{r_{i}}(t)}{\sqrt{p_{r_{i}}}}$:
$\int_{t}[x_{r_{j}}(t)]^{H}\frac{x_{r_{i}}(t)}{\sqrt{p_{r_{i}}}}\ dt=\left\{\begin{aligned} \sqrt{p_{r_{i}}}\ \ \ &j=i,\\ 0\ \ \ &j\neq i,\end{aligned}\right.$ (3)
where $p_{r_{i}}$ is the transmit power of radar $i$. After going through the MF and being sampled by the analog-to-digital converter, the expression of the discrete echo is given by [23]
$y_{r_{i}}(n)$ $=[h_{{r_{i}},i}(n)]^{H}\sqrt{p_{r_{i}}(n)}+[\mathbf{h}_{c\rightarrow r_{i}}(n)]^{H}\sqrt{\mathbf{P}_{c}(n)}\mathbf{e}(n)$ (4)
$+w_{r_{i}}(n),n=1,...,N,$
where $n$ is the discrete time index and $N$ is the sampling number in a coherent processing interval (CPI). For the sake of convenient expression, we omit the index $n$ in the following expression. $h_{r_{i},i}$ is the $i$th element of $\mathbf{h}_{r_{i}}$, which is rightly the two-way link between radar $i$ and the target. $\mathbf{P}_{c}=\text{diag}\{p_{c_{1}},...,p_{c_{M_{c}}}\}$ is the transmit power matrix of BSs. Similarly, we denote the transmit power matrix of radars as $\mathbf{P}_{r}=\text{diag}\{p_{r_{1}},...,p_{r_{M_{r}}}\}$. $w_{r_{i}}$ is the additive white Gaussian noise, which conforms to the $\mathcal{CN}(0,\sigma^{2})$ distribution and $\mathbf{e}\in\mathbb{C}^{M_{c}\times 1}$ is the MF output of communication signals, which conforms to the $\mathcal{CN}(0,\mathbf{I}_{M_{c}})$ distribution. Based on (4), we calculate the received SINR of radar $i$, which is a function of the transmit power of BSs and radars, i.e., $\rho_{i}(\mathbf{P}_{c},\mathbf{P}_{r})$:
@@
$=\frac{[h_{r_{i},i}]^{H}p_{r_{i}}h_{r_{i},i}}{\mathbb{E}\big{(}\mathbf{e}^{H}\sqrt{\mathbf{P}_{c}}\mathbf{h}_{c\rightarrow r_{i}}[\mathbf{h}_{c\rightarrow r_{i}}]^{H}\sqrt{\mathbf{P}_{c}}\mathbf{e}+[w_{r_{i}}]^{H}w_{r_{i}}\big{)}}$
$=\frac{[h_{r_{i},i}]^{H}p_{r_{i}}h_{r_{i},i}}{\sum\limits_{j=1}^{M_{c}}l_{c_{j}\rightarrow r_{i}}p_{c_{j}}+\sigma^{2}}$
$\triangleq\frac{[h_{r_{i},i}]^{H}p_{r_{i}}h_{r_{i},i}}{\sigma_{r}^{2}(\mathbf{P}_{c})},$
@@
where $\sigma_{r}^{2}(\mathbf{P}_{c})$ denotes the interference-plus-noise suffered by radar $i$. The radar uses the generalized likelihood ratio test to judge the presence of the target. The detection probability and false alarm probability, denoted as $P_{D_{i}}$ and $P_{F_{i}}$, are calculated as follows [15],
@@
In terms of the communication system, the received signal of the interfered user is given by
$\mathbf{y}_{c}=[\mathbf{H}_{c}]^{H}\mathbf{x}_{c}+\sum_{i=1}^{M_{r}}[\mathbf{h}_{r_{i}\rightarrow c}]^{H}x_{r_{i}}+\mathbf{w}_{c},$ (7)
where $\mathbf{x}_{c}\in\mathbb{C}^{M_{c}\times 1}=[x_{c_{1}},...,x_{c_{Mc}}]^{T}$ and $\mathbf{x}_{r}=[x_{r_{1}},...,x_{r_{M_{r}}}]^{T}\in\mathbb{C}^{M_{r}\times 1}$ are the transmit signals of BSs and radars, and $\mathbf{w}_{c}\in\mathbb{C}^{N_{c}\times 1}$ is the white noise that conforms to the $\mathcal{CN}(0,\sigma^{2}\mathbf{I}_{N_{c}})$ distribution. Given the large-scale CSI, the ergodic rate of the interfered user, denoted as $\bar{R}(\mathbf{P}_{c},\mathbf{P}_{r})$, can be calculated:
$\bar{R}(\mathbf{P}_{c},\mathbf{P}_{r})=\mathbb{E}_{\mathbf{S}_{c}}\big{[}\log_{2}\text{det}(\mathbf{I}_{N_{c}}+\frac{[\mathbf{H}_{c}]^{H}\mathbf{P}_{c}\mathbf{H}_{c}}{\sigma_{c}^{2}(\mathbf{P}_{r})}\big{)}],$ (8)
where
$\sigma_{c}^{2}(\mathbf{P}_{r})\triangleq\sum\limits_{i=1}^{M_{r}}l_{r_{i}\rightarrow c}p_{r_{j}}+\sigma^{2},$ (9)
which denotes the interference-plus-noise suffered by the user.

### v2

@@ II-B CHANNEL MODEL
In this subsection, we introduce the channel model. We assume that BSs and radars are equipped with a single antenna. This is because, in MIMO systems, the precoding scheme that exploits the multiplexing gain requires perfect and instantaneous CSI [50]. With only large-scale CSI, multiple antennas are equivalent to a combined antenna for both BSs and radars. The interfered user is assumed to have $N_{c}$ antennas, which are used to receive signals from distributed BSs. On this basis, we denote the BS-user link and interference links, i.e., BS-radar link and radar-user link as $\mathbf{H}_{c}^{H}=[\mathbf{h}_{c_{1}}^{H},\mathbf{h}_{c_{2}}^{H},...,\mathbf{h}_{c_{M_{c}}}^{H}]\in\mathbb{C}^{N_{c}\times M_{c}}$, $\mathbf{h}_{c\rightarrow r_{i}}\in\mathbb{C}^{M_{c}\times 1}$, and $\mathbf{h}_{r_{i}\rightarrow c}\in\mathbb{C}^{1\times N_{c}}$, respectively, where $i$ denotes the index number of radars. They are modeled by the composite channel model:
@@
where $\mathbf{L}_{c}\in\mathbb{C}^{M_{c}\times M_{c}}$, $\mathbf{L}_{c\rightarrow r_{i}}\in\mathbb{C}^{M_{c}\times M_{c}}$, and $l_{r_{i}\rightarrow c}$ denote the large-scale fading. $\mathbf{L}_{c}$ and $\mathbf{L}_{c\rightarrow r_{i}}$ are diagonal matrices whose $j$-th diagonal elements are denoted as $l_{c_{j}}$ and $l_{c_{j}\rightarrow r_{i}}$. $\mathbf{S}_{c}\in\mathbb{C}^{M_{c}\times N_{c}}$, $\mathbf{s}_{c\rightarrow r_{i}}\in\mathbb{C}^{M_{c}\times 1}$, and $\mathbf{s}_{r_{i}\rightarrow c}\in\mathbb{C}^{1\times N_{c}}$ denote the small-scale fading, whose elements are uniformly and independently conformed to the Gaussian distribution, i.e., $\mathcal{CN}(0,1)$. For the two-way links between the radars and target, i.e., $\mathbf{H}_{r}=[\mathbf{h}_{r_{1}},....,\mathbf{h}_{r_{M_{r}}}]\in\mathbb{C}^{M_{r}\times M_{r}}$, they are modeled by the determined channel model. This is because when we consider the target in the sky, the corresponding air-to-ground channel is dominated by the line-of-sight path with little random fading.
@@ II-C SIGNAL MODEL
$y_{r_{i}}(t)=$ $\sum_{j=1}^{M_{r}}h^{H}_{r_{j}\rightarrow r_{i}}(t)\sqrt{p_{r_{j}}}\bigg{(}\sum_{n=0}^{N-1}u_{r_{j}}(t-\tau_{r_{j}\rightarrow r_{i}}-n\tau)\bigg{)}e^{j2\pi(f_{c}+f_{d_{r_{j}\rightarrow r_{i}}})(t-\tau_{r_{j}\rightarrow r_{i}})}$ (2)
$+$ $\sum\limits_{j=1}^{M_{c}}h^{H}_{c_{j}\rightarrow r_{i}}(t)\sqrt{p_{c_{j}}}\bigg{(}\sum\limits_{n=0}^{N-1}u_{c_{j}}(t-\tau_{c_{j}\rightarrow r_{i}}-n\tau)\bigg{)}e^{j2\pi f_{c}(t-\tau_{c_{j}\rightarrow r_{i}})}+w_{r_{i}}(t),$
$y_{r_{i}}(t)=$ $\alpha_{i,i}h^{H}_{r_{i}\rightarrow r_{i}}(t)\sqrt{p_{r_{i}}}{\chi}_{i,i}(t-\tau_{r_{i}\rightarrow r_{i}}-n\tau,\Delta^{f_{d}}_{r_{i}\rightarrow r_{i}})e^{j2\pi\tilde{f}_{{d_{r_{i}\rightarrow r_{i}}}}t}$ (3)
$+$ $\sum_{j=1,j\neq i}^{M_{r}}\alpha_{j,i}h^{H}_{r_{j}\rightarrow r_{i}}(t)\sqrt{p_{r_{j}}}{\chi}_{j,i}(t-\tau_{r_{j}\rightarrow r_{i}}-n\tau,\Delta^{f_{d}}_{r_{j}\rightarrow r_{i}})e^{j2\pi\tilde{f}_{{d_{r_{j}\rightarrow r_{i}}}}t}$
$+$ $\sum_{j=1}^{M_{c}}h^{H}_{c_{j}\rightarrow r_{i}}(t)\sqrt{p_{c_{j}}}e_{j}(t-\tau_{c_{j}\rightarrow r_{i}}-n\tau)+w_{r_{i}}(t),$
$y_{r_{i}}(n)=$ $\alpha_{i,i}h^{H}_{r_{i}\rightarrow r_{i}}(t^{r}_{i,n})\sqrt{p_{r_{i}}}{\chi}_{i,i}(\Delta^{\tau}_{r_{i}\rightarrow r_{i}},\Delta^{f_{d}}_{r_{i}\rightarrow r_{i}})e^{j2\pi\tilde{f}_{{d_{r_{i}\rightarrow r_{i}}}}t^{r}_{i,n}}$ (4)
$+$ $\sum_{j=1,j\neq i}^{M_{r}}\alpha_{j,i}h^{H}_{r_{j}\rightarrow r_{i}}(t^{r}_{i,n})\sqrt{p_{r_{j}}}{\chi}_{j,i}(\Delta^{\tau}_{r_{j}\rightarrow r_{i}},\Delta^{f_{d}}_{r_{j}\rightarrow r_{i}})e^{j2\pi\tilde{f}_{{d_{r_{j}\rightarrow r_{i}}}}t^{r}_{i,n}}$
$+$ $\sum_{j=1}^{M_{c}}h^{H}_{c_{j}\rightarrow r_{i}}(t^{r}_{i,n})\sqrt{p_{c_{j}}}e_{j}(\Delta^{\tau}_{c_{j}\rightarrow r_{i}})+w_{r_{i}}(t^{r}_{i,n}),n=1,...,N,$
@@
$=\frac{p_{r_{i}}|h_{r_{i}\rightarrow r_{i}}(t^{r}_{i,n})|^{2}|\chi_{i,i}(\Delta^{\tau}_{r_{i}\rightarrow r_{i}},\Delta^{f_{d}}_{r_{i}\rightarrow r_{i}})|^{2}}{\sum\limits_{j=1,j\neq i}^{M_{r}}p_{r_{j}}|h_{r_{j}\rightarrow r_{i}}(t^{r}_{i,n})|^{2}|\chi_{j,i}(\Delta^{\tau}_{r_{j}\rightarrow r_{i}},\Delta^{f_{d}}_{r_{j}\rightarrow r_{i}})|^{2}+\mathbb{E}\bigg{(}\sum\limits_{j=1}^{M_{c}}p_{c_{j}}e_{j}(\Delta^{\tau}_{c_{j}\rightarrow r_{i}})^{H}e_{j}(\Delta^{\tau}_{c_{j}\rightarrow r_{i}})|h_{c_{j}\rightarrow r_{i}}(t^{r}_{i,n})|^{2}+w_{r_{i}}^{H}(t^{r}_{i,n})w_{r_{i}}(t^{r}_{i,n})\bigg{)}}$
$=\frac{p_{r_{i}}|h_{r_{i}\rightarrow r_{i}}|^{2}|\chi_{i,i}(\Delta^{\tau}_{r_{i}\rightarrow r_{i}}\Delta^{f_{d}}_{r_{i}\rightarrow r_{i}})|^{2}}{\sum\limits_{j=1,j\neq i}^{M_{r}}p_{r_{j}}|h_{r_{j}\rightarrow r_{i}}|^{2}|\chi_{j,i}(\Delta^{\tau}_{r_{j}\rightarrow r_{i}},\Delta^{f_{d}}_{r_{j}\rightarrow r_{i}})|^{2}+\sum\limits_{j=1}^{M_{c}}l_{c_{j}\rightarrow r_{i}}p_{c_{j}}+\sigma^{2}},$
Denote the transmit signals of radars and BSs as $\mathbf{x}_{r}(t)=[x_{r_{1}}(t),...,x_{r_{M_{r}}}(t)]^{T}$ and $\mathbf{x}_{c}(t)=[x_{c_{1}}(t),...,x_{c_{M_{c}}}(t)]^{T}$, respectively. The radar emits $N$ pulses within a coherent processing interval (CPI): $N$ repetitions of an orthogonal waveform, i.e., $x_{r_{i}}(t)=\sum\limits_{n=0}^{N-1}\sqrt{p_{r_{i}}}u_{r_{i}}(t-n\tau)e^{j2\pi f_{c}t},\ t\in[0,\tau],n=1,...,N$, where $p_{r_{i}}$ is the transmit power of the $i$-th radar, $u_{r_{i}}(t)$ is the complex envelop of a radar pulse, which has unit energy, $f_{c}$ denotes the carrier frequency, and $\tau$ is the pulse duration. Similarly, the communication signal is expressed as: $x_{c_{i}}(t)=\sum\limits_{n=0}^{N-1}\sqrt{p_{c_{i}}}u_{c_{i}}(t-n\tau)e^{j2\pi f_{c}t},\ t\in[0,\tau],n=1,...,N$, where $p_{c_{i}}$ denotes the transmit power of the $i$-th BS and $\tau$ is the duration of a communication symbol [^2] In this paper, we assume that the pulse duration of radars is the same as the symbol duration of BSs and C&S systems use the same sampling rate, i.e., $\frac{1}{\tau}$ [32][37].. Denote the bistatic delay from the $j$-th radar to the $i$-th radar as $\tau_{r_{j}\rightarrow r_{i}}$, the bistatic delay from the $j$-th BS to the $i$-th radar as $\tau_{c_{j}\rightarrow r_{i}}$, and the Doppler shift from the $j$-th radar to the $i$-th radar as $f_{d_{r_{j}\rightarrow r_{i}}}$. The received signal of the $i$-th radar is given by (2) (shown at the top of the next page), where $w_{r_{i}}(t)$ is the noise received by the $i$-th radar.
The received signal first goes through the down conversion (remove the term $e^{j2\pi f_{c}t}$ in (2)) and then goes through a matched filter (MF) to extract the echoes of its own probing signals. The MF of the $i$-th radar is $u_{i}^{H}(-t)e^{j2\pi\tilde{f}_{{d_{r_{i}\rightarrow r_{i}}}}t}$ [51], where $\tilde{f}_{{d_{r_{i}\rightarrow r_{i}}}}$ is the estimated Doppler Shift. The output signal of the MF is given by (3) (shown at the top of the next page), where $\alpha_{j,i}\triangleq e^{-j2\pi(f_{c}+f_{d_{r_{j}\rightarrow r_{i}}})\tau_{r_{j}\rightarrow r_{i}}}$, $\Delta^{f_{d}}_{r_{j}\rightarrow r_{i}}\triangleq f_{{d_{r_{j}\rightarrow r_{i}}}}-\tilde{f}_{{d_{r_{j}\rightarrow r_{i}}}}$ is the Doppler shift error, $e_{j}(t-\tau_{c_{j}\rightarrow r_{i}}-n\tau)$ is the MF output of the $j$-th BS signal, and ${\chi}_{j,i}(t,f)$ is the ambiguity function, which is given by
[…]
Then, the output signal is sampled at the pulse rate, obtaining the samples at the time instants $t^{r}_{i,n}$, where $t^{r}_{i,n}=\tau_{r_{i}\rightarrow r_{i}}+\Delta^{\tau}_{r_{i}\rightarrow r_{i}}+n\tau,n=1,...,N$, and $\Delta^{\tau}_{r_{i}\rightarrow r_{i}}$ is the delay error. The output samples are given by (4) (shown at the top of the next page), where $\Delta^{\tau}_{r_{j}\rightarrow r_{i}}\triangleq\tau_{r_{i}\rightarrow r_{i}}-\tau_{r_{j}\rightarrow r_{i}}$ and $\Delta^{\tau}_{c_{j}\rightarrow r_{i}}\triangleq\tau_{r_{i}\rightarrow r_{i}}-\tau_{c_{j}\rightarrow r_{i}}$. On this basis, we calculate the averaged SINR of the $i$-th radar, which is given by (5) [^3] We assume the two-way channel between the radar and the target and the large-scale CSI of the interference channels keep unchanged within a CPI, and omit the time index $t^{r}_{i,n}$ in (5). (shown at the top of the next page), where $\mathbf{P}_{r}=\text{diag}\{p_{r_{1}},...,p_{r_{M_{r}}}\}$, $\mathbf{P}_{c}=\text{diag}\{p_{c_{1}},...,p_{c_{M_{c}}}\}$, $e_{j}(\Delta^{\tau}_{c_{j}\rightarrow r_{i}})$ is assumed to conform the $\mathcal{CN}(0,1)$ distribution and $w_{r_{i}}(t^{r}_{i,n})$ conforms the $\mathcal{CN}(0,\sigma^{2})$ distribution.
To focus on the problem, we assume that the delay and Doppler shift are compensated, i.e., $\Delta^{\tau}_{r_{i}\rightarrow r_{i}}=0,\Delta^{f_{d}}_{r_{i}\rightarrow r_{i}}=0$, and the orthogonality of radar waveforms holds across the whole delay-Doppler plane, i.e., ${\chi}_{i,i}(\Delta^{\tau}_{r_{i}\rightarrow r_{i}},\Delta^{f_{d}}_{r_{i}\rightarrow r_{i}})=1$ and ${\chi}_{j,i}(\Delta^{\tau}_{r_{j}\rightarrow r_{i}},\Delta^{f_{d}}_{r_{j}\rightarrow r_{i}})=0$. Under this assumption, the received SINR of the $i$-th radar is given by
$\rho_{i}(\mathbf{P}_{r},\mathbf{P}_{c})=$ (7)
$\frac{p_{r_{i}}|h_{r_{i}\rightarrow r_{i}}|^{2}|\chi_{i,i}(\Delta^{\tau}_{r_{i}\rightarrow r_{i}},\Delta^{f_{d}}_{r_{i}\rightarrow r_{i}})|^{2}}{\sum\limits_{j=1,j\neq i}^{M_{r}}p_{r_{j}}|h_{r_{j}\rightarrow r_{i}}|^{2}|\chi_{j,i}(\Delta^{\tau}_{r_{j}\rightarrow r_{i}},\Delta^{f_{d}}_{r_{j}\rightarrow r_{i}})|^{2}+\sum\limits_{j=1}^{M_{c}}l_{c_{j}\rightarrow r_{i}}p_{c_{j}}+\sigma^{2}}$
$=\frac{p_{r_{i}}|h_{r_{i}\rightarrow r_{i}}|^{2}}{\sum\limits_{j=1}^{M_{c}}l_{c_{j}\rightarrow r_{i}}p_{c_{j}}+\sigma^{2}}\triangleq\frac{p_{r_{i}}|h_{r_{i}\rightarrow r_{i}}|^{2}}{\sigma_{r}^{2}(\mathbf{P}_{c})},$
@@
where $\sigma_{r}^{2}(\mathbf{P}_{c})$ denotes the interference-plus-noise suffered by radar $i$. In our setting, radar uses the generalized likelihood ratio test to judge the presence of the target. The detection probability and false alarm probability, denoted as $P_{D_{i}}$ and $P_{F_{i}}$, are calculated as follows [15],
@@
As for the communication system, since the distributed BSs work in a user-centric manner, they would estimate the transmission delay and compensate delay differences before sending communication signals. Therefore, the interfered user receives nearly aligned signals from distributed BSs, i.e., $\tau_{c}=\tau_{c_{j}},\forall j$:
$\mathbf{y}_{c}(t)=\sum\limits_{j=1}^{M_{c}}\mathbf{h}^{H}_{c_{j}}(t)\bigg{(}\sum\limits_{n=0}^{N-1}\sqrt{p_{c_{j}}}u_{c_{j}}(t-n\tau-\tau_{c})e^{j2\pi f_{c}(t-\tau_{c})}\bigg{)}$ (9)
$+\sum_{j=1}^{M_{r}}\mathbf{h}^{H}_{r_{j}\rightarrow c}(t)\bigg{(}\sum\limits_{n=0}^{N-1}\sqrt{p_{r_{j}}}u_{r_{j}}(t-n\tau-\tau_{r_{j}})e^{j2\pi f_{c}(t-\tau_{r_{j}})}\bigg{)}$
$+\mathbf{w}_{c}(t),$
where we do not consider the Doppler shift by assuming the interfered user is relatively static. Afterwards, the communication signal goes through the down conversion and is sampled at symbol rate: $t^{c}_{n}=n\tau+\tau_{c},n=1,...,N$. The sampled signal is given by
$\mathbf{y}_{c}(n)=\sum\limits_{j=1}^{M_{c}}\mathbf{h}^{H}_{c_{j}}(t^{c}_{n})\sqrt{p_{c_{j}}}u_{c_{j}}(0)e^{j2\pi f_{c}(t^{c}_{n}-\tau_{c})}+$ (10)
$\sum_{j=1}^{M_{r}}\mathbf{h}^{H}_{r_{j}\rightarrow c}(t^{c}_{n})\sqrt{p_{r_{j}}}u_{r_{j}}(\Delta^{\tau}_{r_{j}\rightarrow c})e^{j2\pi f_{c}(t^{c}_{n}-\tau_{r_{j}})}$
$+\mathbf{w}_{c}(t^{c}_{n}),n=1,...,N,$
where $\Delta^{\tau}_{r_{j}\rightarrow c}\triangleq\tau_{c}-\tau_{r_{j}}$ and $\mathbf{w}_{c}(t^{c}_{n})\in\mathbb{C}^{N_{c}\times 1}$ is the white noise that conforms to the $\mathcal{CN}(0,\sigma^{2}\mathbf{I}_{N_{c}})$ distribution. The averaged interference-plus-noise suffered by the interfered user is calculated by
$\sigma_{c}^{2}(\mathbf{P}_{r})=$ $\mathbb{E}\bigg{(}\sum_{j=1}^{M_{r}}p_{r_{j}}\mathbf{h}^{H}_{r_{j}\rightarrow c}(t^{c}_{n})\mathbf{h}_{r_{j}\rightarrow c}(t^{c}_{n})\times$ (11)
$u_{r_{j}}^{H}(\Delta^{\tau}_{r_{j}\rightarrow c})u_{r_{j}}(\Delta^{\tau}_{r_{j}\rightarrow c})+\mathbf{w}_{c}(t^{c}_{n})\mathbf{w}_{c}^{H}(t^{c}_{n})\bigg{)}$
$=$ $\bigg{(}\sum\limits_{j=1}^{M_{r}}l_{r_{j}\rightarrow c}p_{r_{j}}+\sigma^{2}\bigg{)}\mathbf{I}_{N_{c}}.$
Given the large-scale CSI, the ergodic rate of the interfered user, denoted as $\bar{R}(\mathbf{P}_{c},\mathbf{P}_{r})$, can be calculated:
$\bar{R}(\mathbf{P}_{c},\mathbf{P}_{r})=\mathbb{E}_{\mathbf{S}_{c}}\big{[}\log_{2}\text{det}(\mathbf{I}_{N_{c}}+\frac{\mathbf{H}_{c}^{H}\mathbf{P}_{c}\mathbf{H}_{c}}{\sigma_{c}^{2}(\mathbf{P}_{r})}\big{)}].$ (12)
Remark 1:
Since we assume the delay and Doppler shifts are perfectly compensated, (7) and (12) give the optimistic description of the radar SINR and the communication ergodic rate. This aligns to the reality when the delay and Doppler shift errors are small and works effectively for the literature investigation, e.g., [26, 27, 30, 32]. Here, we simply discuss how the delay and Doppler shift are measured and compensated in C&S systems.
As for the radar side, this measurement is implemented by applying a set of MFs with distinct Doppler shifts to match the received echoes and sampling the output at a given rate [52]. Therefore, the delay error and the Doppler shift error are given by: $|\Delta_{\tau}|\leqslant\frac{T_{s}}{2},|\Delta_{f_{d}}|\leqslant\frac{\Delta f_{d}}{2}$, where $T_{s}$ is the sampling frequency and $\Delta_{f_{d}}$ is the frequency interval of the MFs. As for the communication system, the user position is collected for the delay estimation. Based on the estimated delay, BSs compensate the delay differences in advance, ensuring the signals arrive at the same or closely matched time. Then, the receiver further mitigates time discrepancies according to the insert timing signals. The residual tiny differences would not impact the symbol-level decoding. The Doppler shift compensation is achieved by signal processing methods such as the adaptive equalization. In terms of fast-moving scenarios, 5G system uses the extended cycle prefix in OFDM waveforms to bypass the inter-symbol interference caused by the Doppler shift.
Compared with the communication system that serves for the user who is cooperative, the radar needs to detect the unknown target, which makes it more difficult to perfectly compensate the delay and Doppler shift. In the case of no negligible error of the delay and Doppler shift, the auto-term in (5) becomes less than 1 and the cross-term is no longer zero, i.e., $\chi_{i,i}(\Delta^{\tau}_{r_{i}\rightarrow r_{i}}\Delta^{f_{d}}_{r_{i}\rightarrow r_{i}})<1$ and $\chi_{j,i}(\Delta^{\tau}_{r_{j}\rightarrow r_{i}},\Delta^{f_{d}}_{r_{j}\rightarrow r_{i}})>0$, which lowers the value of the received SINR. We discuss the impacts of the inaccurate delay and Doppler shift compensation in the simulation.